A particle is dropped from height 100m and another particle is projected vertically upward with velocity 50m/s from the ground along the same line.find out the position where two particle will meet

Grade:12

2 Answers

Ram agrawal
5 years ago
I'm doing this ques with the concept of relative motion
Let us consider dropped particle A and upward thrown particle B.(suppose u are a insect and sitting on B , now B will be at rest for u)
ย 
Ram agrawal
5 years ago
Now ,
Vabย = Va-Vb (velocity of A w.r.t B)
(while sitting on A seeing B is Vab)
Vabย = 0-50 = -50m/s
ย 
aab=ย  -g-(-g) =0 ( acceleration of A w.r.t B)
Distance=100m (as time and distance between 2 particle are independent of frame)
ย 
Now , since a = 0 it is a simple non-accelerated motion.
ย 
Applying ,
Speed=distance/time
ย 
50=100/t
=>time=2 sec
ย 
ย Now, we can calculate distance travelled by A in time=2sec , that is at which height they will meet.)
ย 
=>S=ut+at2/2ย  ( 2ndย equation of motion)
=>S=(0)(2)-(10)(2)2/2
=>S=-20m ( negative sign signifies direction)
ย 
So, the particles will meet 20m from top or 30m above ground.
ย 
Here is ur requested answer, there is also a suggestion if u understood just do it without seeing above
ย 
ย